# GS-KAN: Parameter-Efficient Kolmogorov-Arnold Networks via Sprecher-Type Shared Basis Functions

### v1

@@ Abstract
The Kolmogorov-Arnold representation theorem offers a theoretical alternative to Multi-Layer Perceptrons (MLPs) by placing learnable univariate functions on edges rather than nodes. While recent implementations such as Kolmogorov-Arnold Networks (KANs) demonstrate high approximation capabilities, they suffer from significant parameter inefficiency due to the requirement of maintaining unique parameterizations for every network edge. In this work, we propose GS-KAN (Generalized Sprecher-KAN), a lightweight architecture inspired by David Sprecher’s refinement of the superposition theorem. GS-KAN constructs unique edge functions by applying learnable linear transformations to a single learnable, shared parent function per layer. We evaluate GS-KAN against existing KAN architectures and MLPs across synthetic function approximation, tabular data regression and image classification tasks. Our results demonstrate that GS-KAN outperforms both MLPs and standard KAN baselines on continuous function approximation tasks while maintaining superior parameter efficiency. Additionally, GS-KAN achieves competitive performance with existing KAN architectures on tabular regression and outperforms MLPs on high-dimensional classification tasks. Crucially, the proposed architecture enables the deployment of KAN-based architectures in high-dimensional regimes under strict parameter constraints, a setting where standard implementations are typically infeasible due to parameter explosion. The source code is available at https://github.com/rambamn48/gs-impl.
1 Introduction
@@
However, standard KANs face a critical scalability bottleneck. The reliance on unique function parameterization for every edge leads to a parameter complexity of $O(C\cdot N_{in}\cdot N_{out})$, where $C$ is the number of parameters for the learned functions. For high-dimensional inputs, such as flattened images, this results in an explosion of parameters that renders standard KANs impractical or impossible to deploy under strict memory budgets. While recent variants like Wav-KAN [4] attempt to address this by substituting splines with fixed wavelets, they often sacrifice the learnable adaptability of the basis function.
In this paper, we revisit the theoretical foundations laid by Sprecher [5] to address this inefficiency. While the original KAN architecture [3] leverages the Kolmogorov-Arnold representation theorem [6], it does not fully exploit Sprecher’s subsequent refinement, which demonstrated that the inner functions can be constructed as transformed versions of a single univariate function. Historically, direct implementations of such “Sprecher networks” were hindered by the pathological, fractal nature of the exact functions required by the theorem, rendering them unsuitable for gradient descent, see work on such networks by [7].
@@
  • Parameter Efficiency: We propose a memory-efficient formulation using layer-wise shared learnable B-splines, adapted via learnable linear parameters. This architecture drastically reduces the parameter footprint compared to standard KANs.
@@
  • Empirical Validation: We demonstrate that GS-KAN is highly versatile. It outperforms MLPs and KAN architectures on high-frequency function approximation and matches or exceeds State-of-the-Art KAN variants on tabular benchmarks. Furthermore, we show that GS-KAN can successfully classify high-dimensional image vectors (Fashion-MNIST) with fewer parameters than MLPs.
@@
2 Theoretical Background
@@
2.1 Universal Approximation Theorem (MLP)
@@
2.2 Kolmogorov-Arnold Representation (KAN)
@@
2.3 Sprecher’s Refinement
Sprecher [5] significantly refined the representation theory by demonstrating that distinct functions for each branch are not required. Instead, the theorem guarantees that a multivariate function can be represented using single continuous functions $\Phi$ and $\psi$ through a specific construction involving scaling and translation.
@@ 2.3 Sprecher’s Refinement
For any continuous function $f:[0,1]^{n}\to\mathbb{R}$, there exist real constants $\lambda,\epsilon$ and a continuous function $\Phi:\mathbb{R}\to\mathbb{R}$ such that:
$f(\mathbf{x})=\sum_{q=0}^{2n}\Phi\left(\sum_{p=1}^{n}\lambda^{p\cdot q}\psi(x_{p}+\epsilon q)\right)$ (3)
Here, $\psi$ is a universal monotonic function shared across all terms. Crucially, the outer function $\Phi$ is also shared (independent of $q$), acting on the aggregated summation.
2.4 Related Efficient Architectures
@@ 3.1 Generalized Representation
where $\lambda_{p,q}$ represents a learnable edge weight and $\epsilon_{q}$ is a learnable node-specific bias (translation).
@@
3.2 Network Architecture & Implementation
@@
    Static Grid Definition: We fix the definition domain of the B-splines to a static interval [G,G][-G,G] (typically [3,3][-3,3]).
@@
3.3 Parameter Complexity Analysis
The primary motivation for GS-KAN is parameter efficiency. Consider a layer with $N$ inputs and $M$ outputs. Let $C$ be the number of parameters per learned function (e.g. B-splines).
  • Standard KAN: Requires a unique function for every edge. Parameter count $\approx N\times M\times C$.
  • MLP: Requires one weight per edge. Parameter count $\approx N\times M$.
  • GS-KAN: Requires one weight matrix ($\lambda$) and one shared spline ($\psi$). Parameter count $\approx N\times M+C$.
Since $C$ typically is negligible compared to the weight matrix ($N\times M$), GS-KAN achieves the same asymptotic parameter complexity as an MLP ($O(N\cdot M)$), while retaining the mathematical properties of Kolmogorov-Arnold networks. This contrasts with Standard KANs, which are fundamentally heavier ($O(C\cdot N\cdot M)$).
4 Experiments
@@
4.1 Synthetic Function Approximation
@@ Results.
The comparison was conducted in two parameter regimes: Nano ($\approx 200$ params) and Micro ($\approx 500$ params). The specific architectures (depth and width) were adjusted to match the parameter counts across models. Results are presented in Table 1.
Table 1: Function Approximation Results. Comparison of the best test MSE (Mean $\pm$ Std). The Architecture column denotes $N_{in},N_{hidden},\dots,N_{out}]$. Resolution refers to Knots ($K$) for GS-KAN and Grid Size ($G$) for Std-KAN.
@@
4.2 Tabular Regression (California Housing)
@@
We benchmarked performance across three parameter regimes. The specific architectures (depth and width) were adjusted to match the parameter counts across models. Results are shown in Table 2.
Table 2: California Housing Results. Comparison of Best Test MSE (Mean $\pm$ Std). The best result in each regime is highlighted in bold. Note that GS-KAN outperforms the MLP baseline in all regimes.
@@
GS-KAN demonstrates robust scaling properties, consistently outperforming the MLP baseline across all tested parameter regimes. While the MLP exhibits saturation, GS-KAN effectively captures high-frequency interactions. Moreover, GS-KAN remains highly competitive with state-of-the-art variants like Std-KAN and Wav-KAN, frequently matching or exceeding their performance. For instance, in the parameter-constrained Small regime GS-KAN achieves an MSE of $0.236$, surpassing Wav-KAN ($0.244$) and outperforms Std-KAN in the Nano regime. This suggests that the shared-basis architecture offers a highly parameter-efficient alternative.
4.3 High-Dimensional Classification (Fashion-MNIST)
@@
Table 3: Classification Results (Fashion-MNIST). Comparison of Best Test Accuracy with a restricted parameter budget ($\approx 12.5k$). The MLP baseline was granted a larger parameter count to test the efficiency of GS-KAN.
@@
5 Conclusion and Future Work
@@
    Competitive Real-World Performance: On tabular benchmarks, GS-KAN matches the performance of State-of-the-Art KAN variants, often yielding superior accuracy-per-parameter ratios in low-resource settings.

### v2

@@ Abstract
The Kolmogorov-Arnold representation theorem offers a theoretical alternative to Multi-Layer Perceptrons (MLPs) by placing learnable univariate functions on edges rather than nodes. While recent implementations such as Kolmogorov-Arnold Networks (KANs) demonstrate high approximation capabilities, they suffer from significant parameter inefficiency due to their reliance on unique parameterizations for every network edge. In this work, we propose GS-KAN (Generalized Sprecher-KAN), a lightweight architecture inspired by David Sprecher’s refinement of the superposition theorem. GS-KAN constructs unique edge functions by applying learnable linear transformations to a single learnable, shared parent function per layer. We evaluate GS-KAN against existing KAN architectures and MLPs across synthetic function approximation, tabular data regression, and image classification tasks. Our results demonstrate that GS-KAN outperforms both MLPs and standard KAN baselines on continuous function approximation while maintaining superior parameter efficiency. Additionally, GS-KAN achieves competitive performance with existing KAN architectures on tabular regression and outperforms MLPs on high-dimensional classification tasks. Crucially, the proposed architecture enables the deployment of KAN-based architectures in high-dimensional regimes under strict parameter constraints, a setting where standard implementations are typically infeasible due to parameter explosion. The source code is available at https://github.com/rambamn48/gs-impl.
I Introduction
@@
However, standard KANs face a critical scalability bottleneck. The reliance on unique function parameterization for every edge leads to a parameter complexity of $O(C\cdot N_{in}\cdot N_{out})$, where $C$ is the number of parameters for the learned functions. For high-dimensional inputs, such as flattened images, this results in an explosion of parameters that renders standard KANs impractical or prohibitive to deploy under strict memory budgets. While recent variants like Wav-KAN [4] attempt to address this by substituting splines with fixed wavelets, they often sacrifice the learnable adaptability of the basis function.
In this paper, we revisit the theoretical foundations laid by Sprecher [5] to address this inefficiency. While the original KAN architecture [3] leverages the Kolmogorov-Arnold representation theorem [6], it does not fully exploit Sprecher’s subsequent refinement, which demonstrated that the inner functions can be constructed as transformed versions of a single univariate function. Historically, direct implementations of such “Sprecher networks” were hindered by the pathological, fractal nature of the exact functions required by the theorem, rendering them unsuitable for gradient descent (see, e.g., [7]).
@@
  • Parameter Efficiency: We propose a memory-efficient formulation using layer-wise shared learnable B-splines, adapted via learnable linear parameters. This architecture significantly reduces the parameter footprint compared to standard KANs.
@@
  • Empirical Validation: We demonstrate that GS-KAN is highly versatile. It outperforms MLPs and KAN architectures on high-frequency function approximation and matches or exceeds existing KAN variants on tabular benchmarks. Furthermore, we show that GS-KAN can successfully classify high-dimensional image vectors (Fashion-MNIST) with fewer parameters than MLPs.
A recent concurrent study by [8] also explores the application of Sprecher’s theorem for parameter efficiency. Their approach focuses on a strict realization of Sprecher’s specific summation formula using shared weight vectors. In contrast our work builds upon a variation of the theorem and from this basis, we propose a generalized architecture where we relax the fixed constant constraints to construct unique edge functions via learnable per-edge linear transformations. This distinction allows our method to retain the memory benefits of the theorem while adopting the flexible connectivity typical of modern deep learning layers.
@@
[…]
Figure 1: Edge-Node Computation Comparison. Left: Standard KAN learns a unique function $\phi_{q,p}$ for every connection. Right: GS-KAN learns a single shared basis $\psi$ and adapts it to each edge via learnable scalars $\lambda$ and shifts $\epsilon$. The node $y_{q}$ aggregates the results.
II Theoretical Background
@@
II-A Universal Approximation Theorem (MLP)
@@
II-B Kolmogorov-Arnold Representation (KAN)
@@
II-C Sprecher’s Refinement
Sprecher [5] significantly refined the representation theory by demonstrating that distinct functions for each branch are not required. Instead, the theorem guarantees that a multivariate function can be represented using single continuous functions $\chi$ and $\psi$ through a specific construction involving scaling and translation.
@@ II-C Sprecher’s Refinement
For any continuous function $f:[0,1]^{n}\to\mathbb{R}$, there exist real constants $\lambda,\epsilon$ and a continuous function $\chi:\mathbb{R}\to\mathbb{R}$ such that:
$f(\mathbf{x})=\sum_{q=0}^{2n}\chi\left(\sum_{p=1}^{n}\lambda^{p\cdot q}\psi(x_{p}+\epsilon q)\right)$ (3)
Here, $\psi$ is a universal function shared across all terms. Crucially, the outer function $\chi$ is also shared (independent of $q$), acting on the aggregated summation.
II-D Related Efficient Architectures
@@ III-A Generalized Representation
where $\lambda_{p,q}$ represents a learnable edge weight and $\epsilon_{q}$ is a learnable node-specific bias (translation). See Figure 1 for an illustration of this construction.
@@
III-B Network Architecture & Implementation
@@
    Static Grid Definition: We fix the definition domain of the B-splines to a static interval [G,G][-G,G].
@@
III-C Parameter Complexity Analysis
The primary motivation for GS-KAN is parameter efficiency. Consider a layer with $N_{in}$ inputs and $N_{out}$ outputs. Let $C$ be the number of parameters per learned function (e.g., B-splines coefficients). For simplicity, we omit linear bias/shift terms ($N_{out}$) for both MLP and GS-KAN in this comparison.
  • Standard KAN: Requires a unique function for every edge. Parameter count $\approx N_{in}\times N_{out}\times C$.
  • MLP: Requires one weight per edge. Parameter count $\approx N_{in}\times N_{out}$.
  • GS-KAN: Requires one weight matrix ($\lambda$) and one shared spline ($\psi$). Parameter count $\approx N_{in}\times N_{out}+C$.
Since $C$ typically is negligible compared to the weight matrix ($N_{in}\times N_{out}$), GS-KAN achieves the same asymptotic parameter complexity as an MLP ($O(N_{in}\cdot N_{out})$), while retaining the mathematical properties of Kolmogorov-Arnold networks. This contrasts with Standard KANs, which are fundamentally heavier ($O(C\cdot N_{in}\cdot N_{out})$).
IV Experiments
@@
IV-A Synthetic Function Approximation
@@ Results.
The comparison was conducted in two parameter regimes: Nano ($\approx 200$ params) and Micro ($\approx 500$ params). The specific architectures (depth and width) were adjusted to match the parameter counts across models. Results are presented in Table I and an illustration of the GS-KAN (micro) prediction is found in Figure 2.
Figure 2: Function Approximation Visualization. Left: The ground truth function $f(x,y)=\sin(3\pi x)\cdot\cos(3\pi y)$ (noiseless). Right: The approximation generated by GS-KAN (Micro configuration, $\approx 500$ parameters). Notably, the model effectively captures high-frequency structures and sharp peaks despite strict parameter constraints, a regime where MLPs typically suffer from over-smoothing.
TABLE I: Function Approximation Results. Comparison of the best test MSE (Mean $\pm$ Std). The Architecture column denotes $N_{in},N_{hidden},\dots,N_{out}]$. Resolution refers to Knots ($K$) for GS-KAN and Grid Size ($G$) for Std-KAN.
@@
IV-B Tabular Regression (California Housing)
@@
We benchmarked performance across three parameter regimes. The specific architectures (depth and width) were adjusted to match the parameter counts across models. Results are shown in Table II.
TABLE II: California Housing Results. Comparison of Best Test MSE (Mean $\pm$ Std). The best result in each regime is highlighted in bold. Note that GS-KAN outperforms the MLP baseline in all regimes.
@@
GS-KAN demonstrates robust scaling properties, consistently outperforming the MLP baseline across all tested parameter regimes. While the MLP exhibits saturation, GS-KAN effectively captures high-frequency interactions. Moreover, GS-KAN remains highly competitive with state-of-the-art variants like Std-KAN and Wav-KAN, frequently matching or exceeding their performance. For instance, in the parameter-constrained Small regime GS-KAN achieves an MSE of $0.236$, surpassing Wav-KAN ($0.244$) and outperforms Std-KAN in the Nano regime.
IV-C High-Dimensional Classification (Fashion-MNIST)
@@
TABLE III: Classification Results (Fashion-MNIST). Comparison of Best Test Accuracy with a restricted parameter budget ($\approx 12.5k$). The MLP baseline was granted a larger parameter count to test the efficiency of GS-KAN.
@@
V Conclusion and Future Work
@@
    Competitive Real-World Performance: On tabular benchmarks, GS-KAN matches the performance of current KAN variants, often yielding superior accuracy-per-parameter ratios in low-resource settings.